Write the successor of:
(a) 2440701
(b) 100199
(c) 1099999
(d) 2345670

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(a) Successor of 244070 is 244070 + 1 = 244071 Hence, successor of 244070 is 244071. (b) Successor of 100199 is 100199 + 1 = 100200 Hence, successor of 100199 is 100200. (c) Successor of 1099999 is 1099999 + 1 = 1100000 Hence, successor of 1099999 is 1100000. (d) Successor of 2345670 is 2345670 + 1 = 2345671 Hence, successor of 2345670 is 2345671